OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (anonymous):

You can form the equation (y - 5) = 2(x + 1) from the data you're given by substituting the values into \[y - y_0 = m(x - x_0)\] Then you can manipulate the equation into standard form by removing all fractions via multiplication amd then rearranging to the form Ax + By = C, where A is positive.

OpenStudy (anonymous):

okay